A Has One Share In Which There Is One Priza And Two Blank;B Has Three Shares In Which There Are 3 Prizes And 6 Blanks;Compare The Probability Of A's Success To B's

Grade:12th pass

1 Answers

Arun
4 years ago
W - Winning ticket L - loosing ticket 

WLL, WWL, WWW 
WLL = 3C1 * 9C2 
WWL = 3C2 * 9C1 
WWW = 3C3 * 9C0 

No of ways to chose 3 tickes out of 12 = 12C3 

P(A) = ( 3C1 * 9C2 + 3C2 * 9C1 + 3C3 * 9C0 ) / 12C3 = 34/55 

P(B) = ( 2C2 * 6C0 + 2C1 * 6C1 ) / 8C2 = 12/28